# Traceable Black-Box Watermarks For Federated Learning

### v3

@@ 5.2 Empirical Results
Figure 3: Impact of warmup training ratio $\alpha$ on the MA and VR of $\mathbf{TraMark}$ under IID and non-IID settings.
In the following, we study how each hyperparameter functions of $\mathbf{TraMark}$ in detail.
Warmup Training Ratio $\alpha$.
@@
$\mathbf{TraMark}$ leverages warmup training to achieve better partitioning between the main task region and the watermarking region. Figure 3 presents the averaged MA and VR of $\mathbf{TraMark}$ on four datasets under both IID and non-IID settings with varying warmup training ratios $\alpha$. Notably, $\mathbf{TraMark}$ consistently achieves satisfactory VR with $\alpha\leq 0.5$. When $\alpha=0.7$, $\mathbf{TraMark}$ fails to inject effective watermarks into each personalized global model on time, leading to degraded VR. For both settings, a larger $\alpha$ generally results in a higher MA. Specifically, under the non-IID setting, $\mathbf{TraMark}$ with the default $\alpha=0.5$ achieves an average MA of $64.15\%$, which is only $0.55\%$ lower than FedAvg but $4.65\%$ higher than $\mathbf{TraMark}$ without warmup training. These results highlight the importance of warmup training, as it enables $\mathbf{TraMark}$ to accurately assign unimportant parameters to the watermarking region, thereby minimizing the negative impact of watermarking on main task performance in watermarked models.
Partition Ratio $k$.
@@
The partition ratio $k$ controls the size of the watermarking region. Intuitively, a small $k$ may hinder the watermark injection process as the watermarking region is unable to learn watermark-related information completely. We vary the partition ratio $k$ from $0.1\%$ to $5\%$ to examine its impact on the performance of $\mathbf{TraMark}$. The averaged MA and VR results across all datasets are shown in the left sub-figure of Figure 4. As expected, a smaller $k$ leads to a significant drop in VR, while MA remains nearly unchanged. For example, compared to $\mathbf{TraMark}$ with the default setting ($k=1.0\%$), reducing $k$ to $0.5\%$ causes VR to drop from $99.17\%$ to $84.17\%$, whereas MA shows only a slight increase from $65.66\%$ to $65.70\%$. Moreover, with an extreme value of $k=5.0\%$, MA only drops to $65.16\%$, resulting in a gap of less than $1\%$, while achieving full VR. Therefore, selecting an appropriate $k$ requires balancing MA and VR, with $k=1.0\%$ serving as a practical choice that ensures both reliable watermark injection and minimal performance degradation.
Size of Watermark Dataset.
@@
A larger watermark dataset may improve the effectiveness of the watermark injection process by increasing the diversity and representativeness of watermark triggers, allowing the model to learn a more robust mapping between the watermark trigger and its intended response. To validate this, we vary the size of the watermark dataset from $50$ to $200$ samples. The averaged MA and VR results across all datasets are summarized in the right sub-figure of Figure 4. Similar to the effect of the partition ratio $k$, the size of the watermark dataset significantly impacts the traceability of watermarked models, while having minimal effect on MA. Specifically, when the watermark dataset contains only $50$ triggers, $\mathbf{TraMark}$ achieves a suboptimal VR of $54.17\%$, despite obtaining the highest MA of $65.85\%$. However, with $100$ or more samples, $\mathbf{TraMark}$ ensures successful watermark injection, with only a limited MA drop (at most $0.34\%$), striking a better balance between MA and VR. Therefore, choosing a sufficiently large watermark dataset, such as $100$ samples or more, is essential to ensure the effectiveness of $\mathbf{TraMark}$.
Figure 4: Impact of partition ratio $k$ and watermark dataset size on the performance of $\mathbf{TraMark}$.
@@ 6 Conclusion
We formalize the problem of injecting traceable black-box watermarks in FL. We propose $\mathbf{TraMark}$, which creates a personalized, traceable watermarked model for each client. $\mathbf{TraMark}$ first constructs a personalized global model for each client via masked aggregation. Subsequently, the watermarking process is exclusively performed in the watermarking region of each model using a distinct watermark dataset. The personalized watermarked models are then sent back to each client for local training or deployment. Extensive experiments demonstrate the effectiveness of $\mathbf{TraMark}$ in various FL settings. Additionally, we conduct a comprehensive hyperparameter study of $\mathbf{TraMark}$.
Acknowledgments
This material is based upon work supported by the U.S. Department of Energy, Office of Science, Office of Advanced Scientific Computing Research under Contract No. DE-AC05-00OR22725. This manuscript has been co-authored by UT-Battelle, LLC under Contract No. DE-AC05-00OR22725 with the U.S. Department of Energy. The United States Government retains and the publisher, by accepting the article for publication, acknowledges that the United States Government retains a non-exclusive, paid-up, irrevocable, world-wide license to publish or reproduce the published form of this manuscript, or allow others to do so, for United States Government purposes. The Department of Energy will provide public access to these results of federally sponsored research in accordance with the DOE Public Access Plan (http://energy.gov/downloads/doe-public-access-plan).
@@ A.1 Notation Table
Table 2: Notation Table
@@
$\mathrm{Div}(\cdot)$ The divergence measurement function
@@ A.2 Related Work
Protecting the IP of FL models has been extensively studied recently, with most existing approaches leveraging either parameter-based (Uchida et al. 2017; Darvish Rouhani, Chen, and Koushanfar 2019; Chen et al. 2023; Liang and Wang 2023; Li, Wang, and Liew 2022; Zhang et al. 2024; Yu et al. 2023; Xu et al. 2024) or backdoor-based watermarking (Tekgul et al. 2021; Liu et al. 2021; Li et al. 2022; Shao et al. 2024; Luo and Chow 2024; Nie and Lu 2024b, a; Wu et al. 2022). While both approaches aim to verify model ownership, backdoor-based methods are more practical as they do not require access to model parameters. However, ensuring traceability, i.e., identifying the specific source of a leaked model, remains an open challenge for backdoor-based watermarks.
Parameter-based Watermarking.
Parameter-based watermarking methods typically embed cryptographic information directly into the parameter space of the global model. For example, Uchida et al. (2017) proposed the first watermarking method for DNNs by incorporating a regularization loss term to embed a watermark into the model weights. Similarly, FedIPR (Li et al. 2022) embeds messages in the Batch Normalization layers by assigning each client a random secret matrix and a designated embedding location. However, during verification, the verifier must access the model parameters to extract the embedded information. Consequently, these approaches assume that the verifier has full access to the suspect model, which is often unrealistic in real-world scenarios where leaked models may be only partially accessible (e.g., via API queries) (Lansari et al. 2023).
Table 3: Comparison of watermarking methods in FL. We evaluate whether each method supports black-box verification ($B$), operates entirely on the server side ($S$), and enables traceability of the local model (Traceable). Our method, $\mathbf{TraMark}$, is the only approach that satisfies all three properties.
@@
WAFFLE (Tekgul et al. 2021)
FedIPR (Li et al. 2022)
FedTracker (Shao et al. 2024)
RobWe (Xu et al. 2024)
FedCRMW (Nie and Lu 2024a)
@@
Backdoor-based Watermarking.
Backdoor-based watermarking has been explored as a more practical alternative, as it does not require direct access to the model’s internal parameters. These methods leverage backdoor injection techniques to ensure that the model learns a specific trigger. A watermarked model outputs predefined responses when presented with inputs containing the trigger (Adi et al. 2018). For instance, WAFFLE (Tekgul et al. 2021) generates a global trigger dataset and fine-tunes the global model on it in each training round, thereby embedding the trigger into the model. Similarly, Liu et al. (2021) assume the presence of an honest client in the system and injects a trigger set (constructed by sampling Gaussian noise) through local training. While these methods enable black-box verification, their watermarked model lacks traceability. Moreover, some approaches rely on client-side trigger injection, which poses a high risk of exposure if a malicious client becomes aware of the process.
Traceability of Watermarked Models.
To ensure the traceability of watermarks, Yu et al. (2023) propose replacing the linear layer of a suspect model with a verification encoder that produces distinct responses if the model originates from the FL system. FedTracker (Shao et al. 2024) extends WAFFLE by injecting a trigger into the global model while embedding local fingerprints (key matrices and bit strings) for individual clients. A recent work, RobWe (Xu et al. 2024), follows a similar workflow to ours, splitting the network into two parts: one for model utility and another for embedding watermarks (key matrices). However, since watermark injection occurs on the client side, this approach is less practical. Another client-side method FedCRMW (Nie and Lu 2024a), proposes a collaborative ownership verification method that indicates the leaker by the consensus of results of multiple watermark datasets. However, the watermark dataset used by FedCRMW is constructed based on the main task dataset, which incurs data privacy risks. We further summarize the characteristics of existing methods in Table 3. Notably, our method, $\mathbf{TraMark}$ , is the only approach that enables fully server-side watermarking with both traceability and black-box verifiability.
@@ A.5 Hardware Settings
All experiments were carried out on a self-managed Linux-based computing cluster running Ubuntu 20.04.6 LTS. The cluster is equipped with eight NVIDIA RTX A6000 GPUs (each with 49 GB of memory) and AMD EPYC 7763 CPUs featuring 64 cores.
@@ A.6 Output Divergence
A.7 Other Sources of Watermark Dataset
Table 4: Generalization of $\mathbf{TraMark}$ across different watermarking datasets. MAs are shown in percentages ($\%$). If a method achieves a full VR, we denote it with “✓”; otherwise, we use “✗”.
Watermark Dataset CIFAR-10 CIFAR-100 Tiny-ImageNet
MA VR MA VR MA VR
MNIST $88.58$ $61.13$ $20.91$
SVHN $88.52_{\geq 0.05}$ $60.92_{\geq 0.05}$ $20.22_{\geq 0.05}$
WafflePattern $88.84_{\geq 0.05}$ $61.31_{\geq 0.05}$ $20.91_{\geq 0.05}$
We also use the SVHN (Netzer et al. 2011) dataset and WafflePattern (Tekgul et al. 2021) as sources for watermark datasets. Notably, WafflePattern consists of images containing only noise with specific patterns. Since SVHN and WafflePattern contain colorful images while FMNIST is grayscale, we conduct experiments on CIFAR-10, CIFAR-100, and Tiny-ImageNet. The MA and VR results are summarized in Table 4. We observe that $\mathbf{TraMark}$ achieves highly similar results regardless of the watermark dataset used. Furthermore, we perform a $\mathrm{T}\text{-}\mathrm{test}$ on each dataset to compare the results of $\mathbf{TraMark}$ using SVHN or WafflePattern against MNIST. The obtained p-values across all main task datasets exceed the commonly used significance threshold of $0.05$, indicating that the differences are not statistically significant. These results demonstrate the generalization ability of $\mathbf{TraMark}$ in selecting different watermarking datasets.
A.8 Large-scale FL with Client Sampling
@@
A.9 More Results on Fine-tuning Attack, Quantization Attack, and Adaptive Attack
More Results on Fine-tuning Attack.
By default, we set the fine-tuning learning rate to $0.01$ (local training learning rate), consistent with the training phase. We gradually increase the fine-tuning learning rate to an extreme value of $0.1$ and perform fine-tuning for $30$ rounds. The resulting MA and VR are shown in Figure 6. Our results demonstrate that $\mathbf{TraMark}$ exhibits strong resilience against fine-tuning attacks: as long as MA remains stable, VR is preserved. When the fine-tuning learning rate becomes excessively large and causes a drop in MA, VR also decreases accordingly. This indicates that the two regions are tightly coupled, making it difficult for malicious clients to remove the embedded watermarks through fine-tuning.
@@ A.9 More Results on Fine-tuning Attack, Quantization Attack, and Adaptive Attack
Quantization Attack.
We assume that malicious clients may quantize their local models to impact the effectiveness of watermarks. Following (Shao et al. 2024), we conduct experiments on watermarked models trained by $\mathbf{TraMark}$ that are quantized to FP16 and INT8. The MA and VR results are summarized in Table 6. Compared to the FP32 baseline, quantizing the model to FP16 and INT8 leads to a $2.09\%$ and $2.10\%$ drop in MA, respectively, and a $0.83\%$ drop in VR. The negligible decrease in VR demonstrates the robustness of the watermarks injected by $\mathbf{TraMark}$ against quantization attacks.
@@
We consider a more challenging scenario where malicious clients are aware that the received global model has been embedded with a black-box watermark. As a result, they attempt to remove the watermark using a backdoor removal method. We adopt Reconstructive Neuron Pruning (RNP) (Li et al. 2023), a state-of-the-art backdoor removal technique, for this purpose. Since RNP requires a Batch normalization layer while ViT employs Layer normalization, we conduct experiments on FMNIST, CIFAR-10, and CIFAR-100. The MA and VR results of RNP on the watermarked models are summarized in Table 7. We observe that RNP has limited effectiveness on the FMNIST and CIFAR-10 datasets, where the VR decreases from $90\%$ and $100\%$ to $70\%$ and $90\%$, respectively. However, in these cases, the MA also drops significantly. For CIFAR-100, RNP has no impact on the VR but still leads to a degradation in MA. These results highlight the robustness of the watermarks embedded in each global model, demonstrating the strong effectiveness of $\mathbf{TraMark}$.
A.10 Extendability of $\mathbf{TraMark}$
As stated in Definition 3 and Remark 1, learning traceable black-box watermarks requires $\phi_{i}(x)\neq\phi_{j}(x),\ \forall i\neq j$ to ensure each watermark dataset produces a distinct output label. Otherwise, the verifier cannot reliably link a model to its corresponding watermark, resulting in failed verification. To meet this condition, we assign a unique label from the main task dataset to each client’s watermark dataset (e.g., client ID $3$ is assigned the one-hot encoding of label $3$ as its output distribution). This naturally limits the number of clients, particularly when the learning task involves only a few labels (e.g., CIFAR-10). To mitigate this limitation, one can extend the model’s output space by introducing synthetic labels—for example, expanding CIFAR-10’s label space from $10$ to $20$ to accommodate $20$ clients. We implement this strategy and observe that $\mathbf{TraMark}$ maintains full VR with only a $0.34\%$ drop in main task accuracy compared to FedAvg, demonstrating its extendability.
A.11 More Discussions and Future Directions
Malicious Client Collusion.

### v4

@@ 5.3 Detailed Studies of $\mathbf{TraMark}$
Figure 3: Impact of warmup training ratio $\alpha$ on the MA and VR of $\mathbf{TraMark}$.
@@
Warmup Training Ratio $\alpha$. $\mathbf{TraMark}$ leverages warmup training to achieve better partitioning between the main task region and the watermarking region. Figure 3 presents the averaged MA and VR of $\mathbf{TraMark}$ on four datasets under both IID and non-IID settings with varying warmup training ratios $\alpha$. Notably, $\mathbf{TraMark}$ consistently achieves satisfactory VR with $\alpha\leq 0.5$. When $\alpha=0.7$, $\mathbf{TraMark}$ fails to inject effective watermarks into each personalized global model on time, leading to degraded VR. For both settings, a larger $\alpha$ generally results in a higher MA. Specifically, under the non-IID setting, $\mathbf{TraMark}$ with the default $\alpha=0.5$ achieves an average MA of $64.15\%$, which is only $0.55\%$ lower than FedAvg but $4.65\%$ higher than $\mathbf{TraMark}$ without warmup training. These results highlight the importance of warmup training, as it enables $\mathbf{TraMark}$ to accurately assign unimportant parameters to the watermarking region, thereby minimizing the negative impact of watermarking on main task performance in watermarked models.
Figure 4: Impact of partition ratio $k$ and watermark dataset size.
@@
Partition Ratio $k$. The partition ratio $k$ controls the size of the watermarking region. Intuitively, a small $k$ may hinder the watermark injection process as the watermarking region is unable to learn watermark-related information completely. We vary the partition ratio $k$ from $0.1\%$ to $5\%$ to examine its impact on the performance of $\mathbf{TraMark}$. The averaged MA and VR results across all datasets are shown in the left sub-figure of Figure 4. As expected, a smaller $k$ leads to a significant drop in VR, while MA remains nearly unchanged. For example, compared to $\mathbf{TraMark}$ with the default setting ($k=1.0\%$), reducing $k$ to $0.5\%$ causes VR to drop from $99.17\%$ to $84.17\%$, whereas MA shows only a slight increase from $65.66\%$ to $65.70\%$. Moreover, with an extreme value of $k=5.0\%$, MA only drops to $65.16\%$, resulting in a gap of less than $1\%$, while achieving full VR. Therefore, selecting an appropriate $k$ requires balancing MA and VR, with $k=1.0\%$ serving as a practical choice that ensures both reliable watermark injection and minimal performance degradation.
@@
Size of Watermark Dataset. A larger watermark dataset may enhance the effectiveness of watermark injection, enabling the model to learn a more robust mapping between the watermark trigger and its intended response. To validate this, we vary the size of the watermark dataset from $50$ to $200$ samples. The averaged MA and VR results across all datasets are summarized in the right sub-figure of Figure 4. Similar to the effect of the partition ratio $k$, the size of the watermark dataset significantly impacts the traceability of watermarked models, while having minimal effect on MA. Specifically, when the watermark dataset contains only $50$ triggers, $\mathbf{TraMark}$ achieves a suboptimal VR of $54.17\%$, despite obtaining the highest MA of $65.85\%$. However, with $100$ or more samples, $\mathbf{TraMark}$ ensures successful watermark injection, with only a limited MA drop (at most $0.34\%$), striking a better balance between MA and VR. Therefore, choosing a sufficiently large watermark dataset, such as $100$ samples or more, is essential to ensure the effectiveness of $\mathbf{TraMark}$.
Table 2: Generalization of $\mathbf{TraMark}$ across different watermarking datasets, including MNIST, SVHN, and WafflePattern. MAs and VRs are shown in percentages ($\%$).
Watermark Dataset CIFAR-10 CIFAR-100 Tiny-ImageNet
MA VR MA VR MA VR
MNIST $88.58$ $100.0$ $61.13$ $100.0$ $20.91$ $100.0$
SVHN $88.52_{p\geq 0.05}$ $100.0$ $60.92_{p\geq 0.05}$ $100.0$ $20.22_{p\geq 0.05}$ $100.0$
WafflePattern $88.84_{p\geq 0.05}$ $100.0$ $61.31_{p\geq 0.05}$ $100.0$ $20.91_{p\geq 0.05}$ $100.0$
Different Type of Watermark Dataset. We also use the SVHN (Netzer et al., 2011) dataset and WafflePattern (Tekgul et al., 2021) as sources for watermark datasets. Since SVHN and WafflePattern contain colorful images while FMNIST is grayscale, we conduct experiments on CIFAR-10, CIFAR-100, and Tiny-ImageNet. The MA and VR results are summarized in Table 2. We observe that $\mathbf{TraMark}$ achieves highly similar results regardless of the watermark dataset used. Furthermore, we perform a $\mathrm{T}\text{-}\mathrm{test}$ on each dataset to compare the results of $\mathbf{TraMark}$ using SVHN or WafflePattern against MNIST. The obtained p-values across all main task datasets exceed the commonly used significance threshold of $0.05$, indicating that the differences are not statistically significant. These results demonstrate the generalization ability of $\mathbf{TraMark}$ in selecting different watermarking datasets. Furthermore, since WafflePattern utilizes synthetic noise with random patterns, it can generate an arbitrary number of unique samples. $\mathbf{TraMark}$ exploits this to assign distinct watermark datasets to every client without the need for real-world data collection.
Table 3: Computational time (seconds) for the watermarking process of WAFFLE, FedTracker, and $\mathbf{TraMark}$.
Method Time (s)
WAFFLE $6.15$
FedTracker $3.63$
FedTracker (per-client) $0.35$
$\mathbf{TraMark}$ $6.85$
$\mathbf{TraMark}$ (per-client) $0.67$
Computational Overhead. We evaluate the server-side computational cost of the watermarking process, with results on CIFAR-10 dataset summarized in Table 3. Since $\mathbf{TraMark}$ generates personalized models to ensure traceability, the computational cost scales linearly with the number of participating clients, a characteristic shared by the state-of-the-art traceable method, FedTracker. However, the absolute overhead remains highly efficient. As shown in the table, the per-client injection time for $\mathbf{TraMark}$ is approximately $0.67$ seconds, which is comparable to FedTracker ($0.35$ seconds). For a system with 10 clients, the total injection time is only $6.85$ seconds per round. Given that cross-silo FL typically involves significant latency from local training and network communication, this server-side overhead is negligible in practice.
Additional Discussions. We have included further analyses to address key aspects of practical deployment. Specifically, Appendix A.9 examines the scalability of $\mathbf{TraMark}$, where new large-scale experiments confirm that our method maintains robust performance as the system scales. In Appendix A.11, we analyze the communication overhead, demonstrating that $\mathbf{TraMark}$ does not impose additional bandwidth costs compared to existing traceable watermarking methods. Appendix A.12 presents a detailed breakdown of per-client MA, confirming that all clients exhibit uniform performance with negligible disparity. Finally, Appendix A.13 discusses the critical trade-off between traceability and client anonymity, arguing that while traceability reduces model-level anonymity, it is a necessary condition for accountability in secure FL settings where the model leaker is a serious concern.
@@ 6 Conclusion
We formalize the problem of injecting traceable black-box watermarks in FL. Based on the problem, we propose $\mathbf{TraMark}$, which creates a personalized, traceable watermarked model for each client. $\mathbf{TraMark}$ first constructs a personalized global model for each client via masked aggregation. Subsequently, the watermarking process is exclusively performed in the watermarking region of each model using a distinct watermark dataset. The personalized watermarked models are then sent back to each client for local training or deployment. Extensive experiments demonstrate the effectiveness of $\mathbf{TraMark}$ across diverse FL settings. We further show that the parameters in the watermarking region are highly coupled with those in the main task region, making the watermarks robust against attacks. We also conduct a comprehensive hyperparameter study of $\mathbf{TraMark}$.
Acknowledgment
This work was supported by the National Science Foundation under Harnessing the Data Revolution for Nevada Fire Science (HDRFS) Seed Grant NSHE-24-37. This material is based upon work co-supported by the U.S. Department of Energy, Office of Science, Office of Advanced Scientific Computing Research under Contract No. DE-AC05-00OR22725. This manuscript has been co-authored by UT-Battelle, LLC under Contract No. DE-AC05-00OR22725 with the U.S. Department of Energy. The United States Government retains and the publisher, by accepting the article for publication, acknowledges that the United States Government retains a non-exclusive, paid-up, irrevocable, world-wide license to publish or reproduce the published form of this manuscript, or allow others to do so, for United States Government purposes. The Department of Energy will provide public access to these results of federally sponsored research in accordance with the DOE Public Access Plan (http://energy.gov/downloads/doe-public-access-plan).
Ethics Statement
This work strictly adheres to the ICLR Code of Ethics.
Reproducibility Statement
We have made every effort to ensure the reproducibility of our results. The implementation of our method is available in a GitHub repository (https://github.com/JiiahaoXU/TraMark).
@@ A.1 Notation Table
Table 4: Notation table.
@@
$\mathtt{Div}(\cdot)$ The divergence measurement function
@@ A.2 Related Work
Protecting the IP of FL models has been extensively studied recently, with most existing approaches leveraging either parameter-based (Uchida et al., 2017; Darvish Rouhani et al., 2019; Chen et al., 2023; Liang and Wang, 2023; Li et al., 2022b; Zhang et al., 2024; Yu et al., 2023; Xu et al., 2024) or backdoor-based watermarking (Tekgul et al., 2021; Liu et al., 2021; Li et al., 2022a; Shao et al., 2024; Luo and Chow, 2024; Nie and Lu, 2024b; a; Wu et al., 2022). While both approaches aim to verify model ownership, backdoor-based methods are more practical as they do not require access to model parameters. However, ensuring traceability, i.e., identifying the specific source of a leaked model, remains an open challenge for backdoor-based watermarks.
Parameter-based Watermarking. Parameter-based watermarking methods typically embed cryptographic information directly into the parameter space of the global model. For example, Uchida et al. (2017) proposed the first watermarking method for DNNs by incorporating a regularization loss term to embed a watermark into the model weights. Similarly, FedIPR (Li et al., 2022a) embeds messages in the Batch Normalization layers by assigning each client a random secret matrix and a designated embedding location. However, during verification, the verifier must access the model parameters to extract the embedded information. Consequently, these approaches assume that the verifier has full access to the suspect model, which is often unrealistic in real-world scenarios where leaked models may be only partially accessible (e.g., via API queries) (Lansari et al., 2023).
Table 5: Comparison of watermarking methods in FL. We evaluate whether each method supports black-box traceability verification ($B$), operates entirely on the server side ($S$), and enables traceability of the local model (Traceable). Our method, $\mathbf{TraMark}$, is the only approach that satisfies all three properties.
@@
WAFFLE (Tekgul et al., 2021)
FedIPR (Li et al., 2022a)
FedTracker (Shao et al., 2024)
RobWe (Xu et al., 2024)
FedCRMW (Nie and Lu, 2024a)
@@
Backdoor-based Watermarking. Backdoor-based watermarking has been explored as a more practical alternative, as it does not require direct access to the model’s internal parameters. These methods leverage backdoor injection techniques to ensure that the model learns a specific trigger. A watermarked model outputs predefined responses when presented with inputs containing the trigger (Adi et al., 2018). For instance, WAFFLE (Tekgul et al., 2021) generates a global trigger dataset and fine-tunes the global model on it in each training round, thereby embedding the trigger into the model. Similarly, Liu et al. (2021) assume the presence of an honest client in the system and injects a trigger set (constructed by sampling Gaussian noise) through local training. While these methods enable black-box verification, their watermarked model lacks traceability. Moreover, some approaches rely on client-side trigger injection, which poses a high risk of exposure if a malicious client becomes aware of the process.
Traceability of Watermarked Models. To ensure the traceability of watermarks, Yu et al. (2023) propose replacing the linear layer of a suspect model with a verification encoder that produces distinct responses if the model originates from the FL system. FedTracker (Shao et al., 2024) extends WAFFLE by injecting a trigger into the global model while embedding local fingerprints (key matrices and bit strings) for individual clients. However, their traceability verification process requires white-box access to the suspicious model’s parameters, which is generally infeasible in practice. A recent work, RobWe (Xu et al., 2024), follows a similar workflow to ours, splitting the network into two parts: one for model utility and another for embedding watermarks (key matrices). However, the model leaker can skip the watermarking step, weaken the embedded watermark, or embed a forged watermark during local training, making the client-side method less practical than the server-side method under our attack model, where all clients are potential model leakers. Another client-side method FedCRMW (Nie and Lu, 2024a), proposes a collaborative ownership verification method that indicates the leaker by the consensus of results of multiple watermark datasets. However, the watermark dataset used by FedCRMW is constructed based on the main task dataset, which incurs data privacy risks. A recent method, MFL-Owner (Gai et al., 2025), targets multi-modal FL by leveraging each client’s visual and language encoders to construct an orthogonal transformation on the client’s trigger set, which serves as a watermark. However, this approach suffers from poor generalizability and fails to scale effectively to broader FL systems. We further summarize the characteristics of existing methods in Table 5.
@@ A.5 Hardware Settings
All experiments were carried out on a self-managed Linux-based computing cluster running Ubuntu 20.04.6 LTS. The cluster is equipped with eight NVIDIA RTX A6000 GPUs (each with 48 GB of memory) and AMD EPYC 7763 CPUs featuring 64 cores.
@@ A.6 Output Divergence
A.7 Large-scale FL with Client Sampling
@@
A.8 More Results on Fine-tuning Attack, Quantization Attack, and Adaptive Attack
@@
Quantization Attack. We assume that malicious clients may quantize their local models to impact the effectiveness of watermarks. Following (Shao et al., 2024), we conduct experiments on watermarked models trained by $\mathbf{TraMark}$ that are quantized to FP16 and INT8. The MA and VR results are summarized in Table 7. Compared to the FP32 baseline, quantizing the model to FP16 and INT8 leads to a $2.09\%$ and $2.10\%$ drop in MA, respectively, and a $0.83\%$ drop in VR. The negligible decrease in VR demonstrates the robustness of the watermarks injected by $\mathbf{TraMark}$ against quantization attacks.
More Results on Fine-tuning Attack. By default, we set the fine-tuning learning rate to $0.01$ (local training learning rate), consistent with the training phase. We gradually increase the fine-tuning learning rate to an extreme value of $0.1$ and perform fine-tuning for $30$ rounds. The resulting MA and VR are shown in Figure 6. Our results demonstrate that $\mathbf{TraMark}$ exhibits strong resilience against fine-tuning attacks: as long as MA remains stable, VR is preserved. When the fine-tuning learning rate becomes excessively large and causes a drop in MA, VR also decreases accordingly. This indicates that the two regions are tightly coupled, making it difficult for malicious clients to remove the embedded watermarks through fine-tuning.
@@
Adaptive Attack. We consider a more challenging scenario where malicious clients are aware that the received global model has been embedded with a black-box watermark. As a result, they attempt to remove the watermark using a backdoor removal method. We adopt Reconstructive Neuron Pruning (RNP) (Li et al., 2023), a state-of-the-art backdoor removal technique, for this purpose. Since RNP requires a Batch normalization layer while ViT employs Layer normalization, we conduct experiments on FMNIST, CIFAR-10, and CIFAR-100. The MA and VR results of RNP on the watermarked models are summarized in Table 8. We observe that RNP has limited effectiveness on the FMNIST and CIFAR-10 datasets, where the VR decreases from $90\%$ and $100\%$ to $70\%$ and $90\%$, respectively. However, in these cases, the MA also drops significantly. For CIFAR-100, RNP has no impact on the VR but still leads to a degradation in MA. These results highlight the robustness of the watermarks embedded in each global model, demonstrating the strong effectiveness of $\mathbf{TraMark}$.
A.9 Scalability Analysis
To evaluate the practical applicability of $\mathbf{TraMark}$ across varying system scales, we discuss its scalability from three critical perspectives: watermark dataset preparation, label space capacity, and computational overhead. We further support this analysis with large-scale experiments on ImageNet-11(Deng et al., 2009).
Watermark Dataset Scalability. Theoretically, $\mathbf{TraMark}$ requires $n$ distinct OOD datasets for $n$ clients to ensure unique traceability. In practice, this requirement poses no barrier to scalability. As demonstrated in our experiments, $\mathbf{TraMark}$ is fully compatible with synthetic pattern generation methods (e.g., WafflePattern). This approach allows for the automated, zero-cost generation of an effectively infinite number of unique trigger sets, thereby eliminating the need to collect or curate real-world OOD data for each new client.
Label Space Scalability via Virtual Class Extension. TraMark requires each client to have a dedicated output slot. Our work focuses on cross-silo FL settings such as collaborations among hospitals or financial institutions, where the number of clients is usually small (for example, <20<20). In such cases, assigning a unique output position to each client is straightforward and adds no structural or operational difficulty. When the number of clients grows beyond the size of the existing output layer, the model can be expanded before FL training begins, for instance, by adding output neurons. We conduct an experiment where the classifier head is enlarged from $10$ to $20$ outputs on CIFAR-10. TraMark maintains full VR with only a $0.34\%$ drop in main accuracy compared to FedAvg, showing that the method remains stable when the output space is extended. This confirms that expanding the label space is a stable operation that does not compromise model utility.
Computational Scalability. The server-side watermark injection process scales linearly with the number of participating clients ($\Theta(n)$). While simulating thousands of clients is computationally intensive, we emphasize that $\mathbf{TraMark}$ is primarily targeted at cross-silo FL settings. In such settings, the linear increase in server-side computation is highly efficient and remains negligible compared to the communication and local training latencies.
Large-scale Validation on ImageNet-11k. To empirically demonstrate scalability beyond standard benchmarks, we conducted a large-scale experiment on the ImageNet-11k dataset with $100$ clients. We utilized the vit-base-patch16-224-in21k (Dosovitskiy et al., 2021) foundation model and WafflePattern triggers. Over 60 training rounds, the system achieved an average MA of $77.80\%$ and a full $100\%$ VR. These results confirm that $\mathbf{TraMark}$ maintains robust performance and traceability even when scaling to larger client pools and complex model architectures.
A.10 More Discussions and Future Directions